OpenStudy (anonymous):

Solve the equation. Type just the numeral answer. Do NOT round your solution. -0.06m = 7.2

OpenStudy (lgbasallote):

so how will you solve 2x = 6?

OpenStudy (lgbasallote):

does my example make sense?

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

no do 6 ÷ 2

OpenStudy (lgbasallote):

nope...when you're given 2x = 6 you divide both sides by 2 2x/2 = 6/2 x = 3

OpenStudy (lgbasallote):

so what do you do in -0.06m = -7.2

OpenStudy (anonymous):

120

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

what about n/-1.5=1.2

OpenStudy (lgbasallote):

what do you think?

OpenStudy (anonymous):

i dont get that one

OpenStudy (lgbasallote):